A table cost 4 times as much as a chair. The table and chair cost $1000 altogether. How much more did the table cost?
Answer by ikleyn(52781)

.

If the chair costs x dollars, then the table cost 4x dollars, according to the condition.


Altogether,


    x + 4x = 1000

    5x     = 1000

     x     = 1000/5 = 200 dollars.


ANSWER.  The table costs 4*200 = 800 dollars.

Solved.